OpenStudy (elisaneedshelp):

Can some1 help me find the height of a cone with a slant height of 8.5 and a radius of 3.2? And then surface area and base area? Please

OpenStudy (anonymous):

 use pythagorean theorem x^2+3.2^2=8.5^2 then to find the base area you would do A=3.2^2 * pi to figure out the surface area the formula is SA=1/2 P(L) + B which stands for P=perimeter (diameter * pi), L= slant height, and add one base area and if you want me to check your answer im willing to do so :) Does that help you a bit?